# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

### 24 Sep 2025 v2

@@ 6.4. An Amplifier of Leadership: Scaling Support, Enabling Growth
In our workshops, participants envisioned how Manager Clone Agents might amplify a manager’s impact by scaling day-to-day guidance to multiple workers at once, widening access to the manager’s attention, and creating space for workers to develop independent judgment. They emphasized that amplification holds potential for helping teams mature, provided that boundaries, escalation paths, and norms against over-reliance are made explicit.
@@
Participants described Manager Clone Agents as useful for coordinating work and giving timely guidance so that tasks do not stall while waiting for a meeting. In W8’s story, the agent greeted him on Monday with a ranked task list, answered common questions, and, when reaching its limits, pinged the manager while suggesting productive interim tasks. W5 similarly imagined an agent from his supervisor accelerating turnaround times, while M1 described lightweight nudges such as a Slack bot that periodically reminds workers of handoffs or check-ins.
@@
Amplification also meant making managers more accessible to workers. W3 explained that a Manager Clone Agent could capture feedback right after a meeting that might otherwise be lost if the real manager was traveling or tied up in back-to-back calls. W4 added that reaching an agent felt more approachable and responsive than sending what she called a “cold email.”
@@
Managers stressed that this kind of amplification works only when agents enforce role boundaries and escalate issues at the right thresholds. M8 envisioned setting clear pre-approved decision bands in her story, noting that “My team can ask my agent if I am okay with certain decisions or then alert me if my authority or insight is needed for a potential project decision. My team would view this as potentially having greater access to me and the resources to get the job done. They would be able to feel more empowered to make decisions and could run thoughts by me more quickly.” At the same time, W10 cautioned that “There’s a difference between giving someone feedback and spoon-feeding them, the latter would create a reliance on constantly having someone to bounce ideas off 24-7.” If the agent is always available, juniors may defer decisions they should learn to own.
@@
Finally, participants described softer forms of amplification, such as carrying culture and care. M5 imagined a Manager Clone Agent that “facilitates focus sessions and mirrors a mentor’s way of thinking on a mentee’s slides or writing”, thereby “passing on both ideas and spirit”. M10 shared how an automated birthday note from a system unexpectedly rekindled a relationship with a former teammate; others saw similar low-stakes, high-touch gestures as well-suited for Manager Clone Agents to maintain warmth and connection across large organizations.
7. Findings: Navigating Risks: Individual, Interpersonal, and Organizational Impacts (RQ2)
@@
While the previous section highlighted how participants envisioned opportunities for extending managerial reach, here we turn to the risks that, in our participants’ imaginaries, could emerge if Manager Clone Agents were poorly integrated into everyday practices. These risks spanned individual, interpersonal, and organizational levels. At the individual level (see subsection 7.1), both managers and workers grappled with shared vulnerabilities around accountability, skill erosion, and replaceability, though their stakes diverged: managers feared threats to authority and identity, while workers worried about blocked career pathways and recognition. At the interpersonal level (see subsection 7.2), trust initially transferred from human relationships to Manager Clone Agent-mediated exchanges, but fragile authenticity, weakened affective cues, and reduced informal contact strained workplace bonds. At the organizational level (see subsection 7.3), efficiency gains were anticipated to flatten hierarchies and accelerate workflows, yet participants feared this would hollow out belonging and weaken culture.
@@ 7.1. Individual Risks: Divided Titles, Shared Vulnerability
Participants often returned to the question of accountability, disagreeing on who carries the blame when a Manager Clone Agent speaks or acts. Concerns centered on errors: agents misinterpreting intent, causing harmful outcomes, or introducing communication mistakes. Managers worried that outputs might drift from their “real” intent because tacit context cannot be cloned, with M10 explaining: “You can’t collect a lifetime of experience into a knowledge base, so some things the agent says won’t be what the leader truly wants”. Others recalled being held responsible for actions they had not authorized in their design fictions, as M6 noted: “You come back and find you’ve carried a lot of blame for things your agent said.” Workers, however, resisted being scapegoated. W2 asked, “If my AI made a decision that later caused problems, is the blame on me, on the AI, or on whom?” and W6 added, “If I revise work per the agent’s advice and it backfires, I’m probably the one who takes the fall.” Participants also anticipated “misunderstandings from transmission errors” (W1) and worried that agents might become a convenient excuse for poor human stewardship: “It gives people something to blame” (W3). As M8 summarized, “If they lean on my avatar and the output becomes distorted, and no one owns it, accountability becomes the issue.”
@@
Participants also framed Manager Clone Agent use in terms of career opportunities, but from opposite vantage points. Some managers such as M1 and M6 imagined reinvesting time saved into managing up and outward: cultivating senior relationships, doing market visits, or expanding scope. Workers in the same workshops described the flip side: sparser access to managers, feeling devalued, and losing motivation. As W2 put it, “I kind of just don’t really feel like working anymore.” For workers, even routine updates or casual chats with supervisors were seen as crucial opportunities to demonstrate impact. W6 explained, “Weekly updates to the boss and some casual chat are themselves an opportunity to build rapport and leave a positive impression. You need to show your manager what impact your work has had, if you want a promotion.” Managers, meanwhile, feared that over-delegation to clone agents would cause skills to atrophy. As M2 warned, “People are greedy, the more they use it, the lazier they become.” To hedge this risk, others stressed the need to preserve core judgment for managers, as M8 insisted: “I want to keep control over my analysis of the metrics. AI can suggest targets, but I still set them.”
@@
Participants further debated the extent to which managerial roles themselves might be replaced. Some workers argued that project managers with limited decision-making authority were highly vulnerable. W6 estimated “I think 80% of his work could really be replaced by an agent.” W2 similarly questioned the uniqueness of managerial labor, remarking: “I don’t think the division of labor made by AI or real managers will be any different, because I feel that neither is perfectly fair.”
@@
Many managers such as M1, M3, M4, M6, M7, M10 and M11 defended higher-order decision-making, creative judgment, and relational work as difficult to automate. Yet some still admitted that agents might still perform adequately to some extent. As M9 reflected, “I would just have to resign to like it just wouldn’t be done maybe the way that I would do it. And that doesn’t mean it’s wrong. It can still do good enough.” Such reflections pushed managers to confront what M9 called “existential and ego connected identity questions,” echoed by M5: “I probably don’t want to authorize him because if my own work could be done by my agent, then what is this job position for me to do?”
@@ 7.2. Interpersonal Risks: Trust Cloned, But Wears with Distance
Trust in Manager Clone Agents initially built on trust in the human managers they represented. Participants explained that because the clone was seen as an extension of a known person rather than an anonymous system, the trust cultivated through prior human relationships could be partially transferred to the agent. In this sense, the “cloned” trust was borrowed credibility: workers felt more comfortable engaging with a Manager Clone Agent when they already trusted the manager behind it. Yet participants also emphasized that this form of trust was fragile. Unlike trust in a human being, which is reinforced by lived interactions, judgment, and emotional reciprocity, cloned trust rested on the assumption that the agent would faithfully mirror the manager’s intent and style. Once communication became mediated by agents, this assumption was tested and often strained by misinterpretations of value, uncertainty about authenticity, and the loss of affective and informal cues.
@@
Participants first highlighted that the credibility of Manager Clone Agents depended on existing human relationships. M12 explained that “people will be more open with Manager Clone Agents than generic agents as it’s talking to someone you know vs. someone you don’t know.” Without this “underlying relationship,” M8 argued, the system would not constitute a “good functioning system.”
@@
Participants then noted that trust weakened when managers relied too heavily on agents for direct interactions. They emphasized that agents could be useful to enhance communication, such as speeding up responses, filling gaps when managers were unavailable, or relaying straightforward information. But when small-group or one-on-one conversations were fully delegated to a clone, the meaning changed. Workers felt devalued, interpreting it as a signal that their work was “unimportant” (W2, W8) or that they were “unworthy of their manager’s time” (W7). This perception, M5 warned, could “diminish motivation.” W7 expressed the frustration more bluntly: “If you can’t be bothered to work with me at all, then why would I respect you?” Managers also worried about this dynamic in reverse, interpreting workers’ reliance on agents as avoidance or disengagement. As M4 explained in another perspective, “Managers wouldn’t like it if workers use a partition layer to block the imagined harm the boss might inflict on them.” In short, participants saw Manager Clone Agents as acceptable supplements to human communication but resisted their use as substitutes for genuine interaction.
@@
Trust was further strained by uncertainty over authenticity. Workers such as W2 and W7 said they often could not tell whether a message came from their manager, from the agent itself, or whether their own contributions were being faithfully conveyed. Because Manager Clone Agents were designed to stand in for real people, this ambiguity directly undermined the very premise of the clone: workers needed to believe it faithfully mirrored their manager, yet could not be sure where the manager ended and the agent began. W10 illustrated this confusion on a storyboard (see Figure 2), showing the frustration that arises when overlapping identities blur accountability and when agents fail to recall past conversations between real individuals.
@@
Figure 2. Story From W10 Showing Frustration When the Manager Clone Agent Fails to Recall Past Conversations.
@@
Participants also emphasized the loss of emotional nuance and informal connection. W2 acknowledged that one benefit was the absence of negative outbursts, since an agent would not lose its temper. Yet even when agents simulated appropriate affect, workers often found it “unconvincing or even demotivating” (W1). The absence of genuine cues reduced moments of casual bonding. As W2 described, “At the beginning of meetings, there is usually a warm-up stage where we chat casually. I think this is when connections between people are built, and that gets lost interacting with agents.” This point sparked debate: M2 felt that authentic bonds were created outside formal settings such as meals or after-hours chats, while W2 considered meeting warm-ups essential. Managers also imagined longer-term erosion of interpersonal ties. M1 wrote in his design story a horrible relationship development that “The number of times I communicate with my staff has decreased, and to me their work is just a schedule sheet on a screen. I can’t even recall the names and appearances of newly hired employees anymore. The anonymous subordinate satisfaction feedback from HR is getting lower and lower.”
@@
Participants stressed the need to balance efficiency with the preservation of interpersonal connection. As M8 remarked, “If I was leading a team, I’d want to make sure that personal connection was still there and that they don’t feel as if they’re just being offloaded.” In their view, trust that is cloned through Manager Clone Agents risks wearing away with distance, gradually reshaping the texture of workplace relationships and, eventually, organizational culture (see subsection 7.3).
@@ 7.3. Organizational Risks: Performance Up, Belonging Down
At the organizational level, participants described a paradox. They anticipated that Manager Clone Agents could improve systemic efficiency: reducing layers of coordination, redistributing managerial responsibilities, and broadening access to leadership. At the same time, they worried that these very efficiencies might erode the social fabric of organizations, weakening belonging and reshaping collective identity. In other words, while Manager Clone Agents promised to streamline structures, participants feared they might hollow out the cultural role that workplaces have long played.
@@
Participants highlighted efficiency as a powerful driver of structural transformation. M10 described how Manager Clone Agents could take over repetitive approvals and routine tasks, relieving managers of “annoyances” and freeing up “an extra hour or two every day.” W6 imagined that if such delegation became common, “maybe there just aren’t that many managers anymore, the organization is becoming increasingly flat.” Echoing this, M10 noted that many intermediary roles such as HR staff or project managers exist mainly “to facilitate communication between managers and front line workers,” positions that agents could absorb. While some participants welcomed this flattening as a path to accessibility and speed, others feared the loss of “competition that historically drove innovation” (M6). Codifying rules into AI also raised new vulnerabilities. W7 warned in his design fiction that malicious workers could exploit system transparency, explaining, “I have now gotten away with embezzlement. I’ll move on to another company in about a year, because of tech turnover, and I’ll get to say I was trusted with a sizeable budget at a major AI-driven company.”
@@
Beyond structure, participants emphasized the risks to organizational and community culture. W1 worried that when most communication flowed through Manager Clone Agents, “it greatly affect this sense of belonging.” M1 dramatized this erosion in a story where “dissatisfied employees were gradually replaced by those who no longer cared”. W2 pushed the scenario further, imagining a future where “The employees of the entire company are all working for AI,” leaving communal identity hollow. M5 satirized the endpoint: “An office filled entirely with mobile AI agents while humans stayed home: some enjoying more family time, others left isolated.”
8. Findings: Rethinking Manager Clone Agent Design in Future Workplace (RQ3)

### 29 Jan 2026 v3

@@ 5.4. An Leadership Amplifier that Scales Mentorship and Allows for Autonomy and Growth
In our workshops, participants envisioned how Manager Clone Agents might amplify a manager’s impact by scaling day-to-day guidance to multiple workers at once, widening access to the manager’s attention, and creating space for workers to develop independent judgment.
@@
Participants described Manager Clone Agents as useful for coordinating work and giving timely guidance so that tasks do not stall while waiting for a meeting. In W15’s story, the agent greeted him on Monday with a ranked task list, answered common questions, and, when reaching its limits, pinged the manager while suggesting productive interim tasks. M9 similarly imagined an agent from his supervisor accelerating turnaround times, while M2 described lightweight nudges such as a Slack bot that periodically reminds workers of handoffs or check-ins.
@@
Amplification also meant making managers more accessible to workers. W7 explained that a Manager Clone Agent could capture feedback right after a meeting that might otherwise be lost if the real manager was traveling or tied up in back-to-back calls. W8 added that reaching an agent felt more approachable and responsive than sending what she called a “cold email.”
@@
Managers stressed that this kind of amplification works only when agents enforce role boundaries and escalate issues at the right thresholds. M16 envisioned setting clear pre-approved decision bands in her story, noting that “My team can ask my agent if I am okay with certain decisions or then alert me if my authority or insight is needed for a potential project decision. My team would view this as potentially having greater access to me and the resources to get the job done. They would be able to feel more empowered to make decisions and could run thoughts by me more quickly.” At the same time, W23 cautioned that “There’s a difference between giving someone feedback and spoon-feeding them, the latter would create a reliance on constantly having someone to bounce ideas off 24-7.” If the agent is always available, juniors may defer decisions they should learn to own.
@@
Finally, participants described softer forms of amplification, such as carrying culture and care. M9 imagined a Manager Clone Agent that “facilitates focus sessions and mirrors a mentor’s way of thinking on a mentee’s slides or writing”, thereby “passing on both ideas and spirit”. M18 shared how an automated birthday note from a system unexpectedly rekindled a relationship with a former teammate; others saw similar low-stakes, high-touch gestures as well-suited for Manager Clone Agents to maintain warmth and connection across large organizations.
6. Findings (RQ2): Navigating Risks: Individual, Interpersonal, and Organizational Impacts
@@
While the previous section highlighted how participants envisioned opportunities for extending managerial reach, here we turn to the risks that, in our participants’ imaginaries, could emerge if Manager Clone Agents were poorly integrated into everyday practices. These risks spanned individual, interpersonal, and organizational levels. At the individual level (see subsection 6.1), both managers and workers grappled with shared vulnerabilities around accountability, skill erosion, and replaceability, though their stakes diverged: managers feared threats to authority and identity, while workers worried about blocked career pathways and recognition. At the interpersonal level (see subsection 6.2), trust initially transferred from human relationships to Manager Clone Agent-mediated exchanges, but fragile authenticity, weakened affective cues, and reduced informal contact strained workplace bonds. At the organizational level (see subsection 6.3), efficiency gains were anticipated to flatten hierarchies and accelerate workflows; yet, participants feared this would hollow out belonging and weaken culture.
@@ 6.1. Individual Risks: Divided Titles, Shared Vulnerability
Participants often returned to the question of accountability, disagreeing on who carries the blame when a Manager Clone Agent speaks or acts. Concerns centered on errors: agents misinterpreting intent, causing harmful outcomes, or introducing communication mistakes. Managers worried that outputs might drift from their “real” intent because tacit context cannot be cloned, with M18 explaining: “You can’t collect a lifetime of experience into a knowledge base, so some things the agent says won’t be what the leader truly wants”. Others recalled being held responsible for actions they had not authorized in their design fictions, as M10 noted: “You come back and find you’ve carried a lot of blame for things your agent said.” Workers, however, resisted being scapegoated. W3 asked, “If my AI made a decision that later caused problems, is the blame on me, on the AI, or on whom?” and W12 added, “If I revise work per the agent’s advice and it backfires, I’m probably the one who takes the fall.” Participants also anticipated “misunderstandings from transmission errors” (W1) and worried that agents might become a convenient excuse for poor human stewardship: “It gives people something to blame” (W7). As M16 summarized, “If they lean on my avatar and the output becomes distorted, and no one owns it, accountability becomes the issue.”
@@
Participants also framed Manager Clone Agent use in terms of career opportunities, but from opposite vantage points. Some managers such as M2 and M10 imagined reinvesting time saved into managing up and outward: cultivating senior relationships, doing market visits, or expanding scope. Workers in the same workshops described the flip side: sparser access to managers, feeling devalued, and losing motivation. As W3 put it, “I kind of just don’t really feel like working anymore.” For workers, even routine updates or casual chats with supervisors were seen as crucial opportunities to demonstrate impact. W12 explained: “Weekly updates to the boss and some casual chat are themselves an opportunity to build rapport and leave a positive impression. You need to show your manager what impact your work has had, if you want a promotion.” Managers, meanwhile, feared that over-delegation to clone agents would cause skills to atrophy. As M4 warned, “People are greedy, the more they use it, the lazier they become.” To hedge this risk, others stressed the need to preserve core judgment for managers, as M16 insisted: “I want to keep control over my analysis of the metrics. AI can suggest targets, but I still set them.”
@@
Participants further debated the extent to which managerial roles themselves might be replaced. Some workers argued that project managers with limited decision-making authority were highly vulnerable. W12 estimated “I think 80% of his work could really be replaced by an agent.” W3 similarly questioned the uniqueness of managerial labor, remarking: “I don’t think the division of labor made by AI or real managers will be any different, because I feel that neither is perfectly fair.”
@@
Many managers such as M2, M5, M6, M10, M13, M18 and M19 defended higher-order decision-making, creative judgment, and relational work as difficult to automate. Yet some still admitted that agents might still perform adequately to some extent. As M17 reflected, “I would just have to resign to like it just wouldn’t be done maybe the way that I would do it. And that doesn’t mean it’s wrong. It can still do good enough.” Such reflections pushed managers to confront what M17 called “existential and ego connected identity questions,” echoed by M9: “I probably don’t want to authorize him because if my own work could be done by my agent, then what is this job position for me to do?”
@@ 6.2. Interpersonal Risks: Trust Cloned, But Wears with Distance
Trust in Manager Clone Agents was initially built on trust in the human managers they represented. Participants explained that because the clone was seen as an extension of a known person rather than an anonymous system, the trust cultivated through prior human relationships could be partially transferred to the agent. In this sense, the “cloned” trust was borrowed credibility: workers felt more comfortable engaging with a Manager Clone Agent when they already trusted the manager behind it. Yet participants also emphasized that this form of trust was fragile. Unlike trust in a human being, which is reinforced by lived interactions, judgment, and emotional reciprocity, cloned trust rested on the assumption that the agent would faithfully mirror the manager’s intent and style. Once communication became mediated by agents, this assumption was tested and often strained by misinterpretations of value, uncertainty about authenticity, and the loss of affective and informal cues.
@@
Participants first highlighted that the credibility of Manager Clone Agents depended on existing human relationships. M21 explained that “people will be more open with Manager Clone Agents than generic agents as it’s talking to someone you know vs. someone you don’t know.” Without this “underlying relationship,” M16 argued, the system would not constitute a “good functioning system.”
@@
Participants then noted that trust weakened when managers relied too heavily on agents for direct interactions. They emphasized that agents could be useful to enhance communication, such as speeding up responses, filling gaps when managers were unavailable, or relaying straightforward information. But when small-group or one-on-one conversations were fully delegated to a clone, the meaning changed. Workers felt devalued, interpreting it as a signal that their work was “unimportant” (W3, W15) or that they were “unworthy of their manager’s time” (W14). This perception, M9 warned, could “diminish motivation.” W14 expressed the frustration more bluntly: “If you can’t be bothered to work with me at all, then why would I respect you?” Managers also worried about this dynamic in reverse, interpreting workers’ reliance on agents as avoidance or disengagement. As M6 explained in another perspective, “Managers wouldn’t like it if workers use a partition layer to block the imagined harm the boss might inflict on them.” In short, participants saw Manager Clone Agents as acceptable supplements to human communication, but resisted their use as substitutes for genuine interaction.
@@
Trust was further strained by uncertainty over authenticity. Workers such as W3 and W14 said they often could not tell whether a message came from their manager, from the agent itself, or whether their own contributions were being faithfully conveyed. Because Manager Clone Agents were designed to stand in for real people, this ambiguity directly undermined the very premise of the clone: workers needed to believe it faithfully mirrored their manager, yet could not be sure where the manager ended and the agent began. W23 illustrated this confusion on a storyboard (see Figure 3), showing the frustration that arises when overlapping identities blur accountability and when agents fail to recall past conversations between real individuals.
@@
Figure 3. Story From W23 Showing Frustration When the Manager Clone Agent Fails to Recall Past Conversations.
A visual related to W23’s Story on Decision-making and administration, highlighting the confusion caused by overlapping identities and the frustration that arises when the agent fails to remember past conversations between real individuals.
@@
Participants also emphasized the loss of emotional nuance and informal connection. W3 acknowledged that one benefit was the absence of negative outbursts, since an agent would not lose its temper. Yet even when agents simulated appropriate affect, workers often found it “unconvincing or even demotivating” (W1). The absence of genuine cues reduced moments of casual bonding. As W3 described, “At the beginning of meetings, there is usually a warm-up stage where we chat casually. I think this is when connections between people are built, and that gets lost interacting with agents.” This point sparked debate: M4 felt that authentic bonds were created outside formal settings such as meals or after-hours chats, while W3 considered meeting warm-ups essential. Managers also imagined longer-term erosion of interpersonal ties. M2 wrote in his design story a horrible relationship development that “The number of times I communicate with my staff has decreased, and to me their work is just a schedule sheet on a screen. I can’t even recall the names and appearances of newly hired employees anymore. The anonymous subordinate satisfaction feedback from HR is getting lower and lower.”
@@
Participants stressed the need to balance efficiency with the preservation of interpersonal connection. As M16 remarked, “If I was leading a team, I’d want to make sure that personal connection was still there and that they don’t feel as if they’re just being offloaded.” In their view, trust that is cloned through Manager Clone Agents risks wearing away with distance, gradually reshaping the texture of workplace relationships and, eventually, organizational culture (see subsection 6.3).
@@ 6.3. Organizational Risks: Performance Up, Belonging Down
At the organizational level, participants described a paradox. They anticipated that Manager Clone Agents could improve systemic efficiency: reducing layers of coordination, redistributing managerial responsibilities, and broadening access to leadership. At the same time, they worried that these very efficiencies might erode the social fabric of organizations, harming inclusiveness, and reshaping collective identity. In other words, while Manager Clone Agents promised to streamline structures, participants feared that they might hollow out the cultural role that workplaces have long played.
@@
The participants highlighted efficiency as a powerful driver of structural transformation. M18 described how Manager Clone Agents could take over repetitive approvals and routine tasks, relieving managers of “annoyances” and freeing up “an extra hour or two every day.” W12 imagined that if such delegation became common, “maybe there just aren’t that many managers anymore, the organization is becoming increasingly flat.” Echoing this, M18 noted that many intermediary roles such as HR staff or project managers exist mainly “to facilitate communication between managers and front line workers,” positions that agents could absorb. While some participants welcomed this flattening as a path to accessibility and speed, others feared: “This will cause the loss of competition that historically drove innovation” (M10).
@@
Apart from the power structure, participants warned about the risks of a fractured culture within the organization community. W1, pointed to inequities that might emerge: “Workers willing to rely heavily on agents could offload routine labor and free up time for other pursuits, while those unwilling or unable to adopt the technology would be forced to carry out more tasks themselves.” She envisioned: “Over time, such asymmetries could create divisions between workers’ groups and amplify perceptions of unfairness.” M2 dramatized this erosion in a story where “dissatisfied employees were gradually replaced by those who no longer cared”. W3 pushed the scenario further, imagining a future where “The employees of the entire company are all working for AI,” leaving communal identity hollow. M9 satirized the endpoint: “An office filled entirely with mobile AI agents while humans stayed home: some enjoying more family time, others left isolated.”
7. Findings (RQ3): Designing for Manager Clone Agents in Future Workplace